Charlieee:

y=log v4(x+5) +5 give the domain and range

jhonyy9:

so when you talk about logarithms what is the domain - where are defined the logarithms ? @Laylalyssa do you know it ?

surjithayer:

\[y=\log \sqrt{4(x+5)}+5\] it is defined if \[4(x+5) \geq 0\] or \[x \ge -5\] so domain is \[x \ge -5\] and range is allreal values